OpenStudy (haleyelizabeth2017):

The function f(x) passes through the point (-3, 3). if g(x)=f(x-3)+4, then which one of these points, if any, must lie on the graph of g(x)? a. (-6, 7) b. (0,7) c. (-6, -1) d. (0, -1) e. none of these

OpenStudy (anonymous):

The prompt gave you a point: \[f \left( -3 \right) = 3\]\[g \left( ? \right) ~~so ~~that ~~f \left( x - 3 \right) = f \left( -3 \right)\]

OpenStudy (anonymous):

Hint: What x-value when added to -3 will give you -3? @haleyelizabeth2017

OpenStudy (haleyelizabeth2017):

0

OpenStudy (anonymous):

Precisely. Now that we know that x = 0, we can cross out A and C.

OpenStudy (haleyelizabeth2017):

Okay

OpenStudy (anonymous):

Can you find anything that you can substitute? BIG HINT: \[~~~~~~~~~~~~~f(−3)=3\]\[~~~~~~~~~~~~~~~~\downarrow\]\[g(0)=f(-3)+4 = ~?\]

OpenStudy (haleyelizabeth2017):

So it is B?

OpenStudy (anonymous):

Yes, exactly. Good job.

OpenStudy (haleyelizabeth2017):

Thank you :)

OpenStudy (anonymous):

You're welcome ^_^